exact (x^2+y^2-5)dx=(y+xy)dy y(0)=1
1
Expert's answer
2021-10-13T12:51:24-0400
"(x^2+y^2-5)dx+(-y-xy)dy=0"


"M(x, y)=x^2+y^2-5, M_y=2y"

"N(x, y)=-y-xy, N_x=-y"

"M_y\not=N_x"

"\dfrac{M_y-N_x}{N}=\dfrac{2y-(-y)}{-y-xy}=-\dfrac{3}{x+1}=P(x)"

Integrating factor


"\mu(x)=e^{\int(-{3 \over x+1})dx}=\dfrac{1}{(x+1)^3}"

"\dfrac{x^2+y^2-5}{(x+1)^3}dx+\dfrac{-y-xy}{(x+1)^3}dy=0"

"M(x, y)=\dfrac{x^2+y^2-5}{(x+1)^3}, M_y=\dfrac{2y}{(x+1)^3}"

"N(x, y)=-\dfrac{y}{(x+1)^2}, N_x=\dfrac{2y}{(x+1)^3}"

"M_y=\dfrac{2y}{(x+1)^3}=N_x"

"f(x, y)=\int(-\dfrac{y}{(x+1)^2})dy+g(x)"

"=-\dfrac{y^2}{2(x+1)^2}+g(x)"

"f_x=\dfrac{y^2}{(x+1)^3}+g'(x)=\dfrac{x^2+y^2-5}{(x+1)^3}"

"g'(x)=\dfrac{x^2-5}{(x+1)^3}"

"g(x)=\int\dfrac{x^2-5}{(x+1)^3}dx=\int\dfrac{x^2+2x+1-2x-2-4}{(x+1)^3}dx"

"=\int\dfrac{1}{x+1}dx-2\int\dfrac{1}{(x+1)^2}dx-4\int\dfrac{1}{(x+1)^3}dx"

"=\ln(|x+1|)+\dfrac{2}{x+1}+\dfrac{2}{(x+1)^2}-C"

"-\dfrac{y^2}{2(x+1)^2}+\ln(|x+1|)+\dfrac{2}{x+1}+\dfrac{2}{(x+1)^2}=C"

The initial condition "y(0)=1"


"-\dfrac{(1)^2}{2(0+1)^2}+\ln(|0+1|)+\dfrac{2}{0+1}+\dfrac{2}{(0+1)^2}=C"

"C=\dfrac{7}{2}"

The solution of the initial value problem


"-\dfrac{y^2}{2(x+1)^2}+\ln(|x+1|)+\dfrac{2}{x+1}+\dfrac{2}{(x+1)^2}=\dfrac{7}{2}"